Three former college classmates have decided to pool a variety of work experiences by opening a store near campus to sell wireless equipment to students. The business has been incorporated as University Wireless.

Required: Several transactions occurred in March. Each is described separately in this folder. For each transaction, indicate the accounts that are affected, whether they increase or decrease, and the amount of the increase or decrease.

YOU MUST FOLLOW THE INSTRUCTIONS BELOW. IF YOU DON'T, YOU MAY KNOW THE CORRECT ENTRY BUT THE COMPUTER WILL NOT RECOGNIZE IT AND YOU WILL NOT RECEIVE CREDIT.

  1. After each transaction description, there are several "Account" submission boxes and corresponding "Amount" submission boxes. To indicate the accounts that you think are affected, choose them from the drop-down menu. But you MUST select them in the order that they are listed in the menu. FOR EXAMPLE, if you think that Cash and Inventory are affected by a particular transaction, you must record the Cash impact first and the Inventory impact second because that is the order in which they are listed in the drop-down menu. If you record the Inventory impact first and the Cash impact second, even if they are the correct accounts and even if you have the correct dollar amounts, your answer will be considered incorrect.
  2. When you record the dollar amounts, be sure to use a minus sign to indicate a decrease in the account. You don't need to use a plus sign to indicate an increase.
  3. There are always more "Account" and "Amount" submission boxes available than are necessary. When you have indicated all the accounts that are affected by the transaction, you MUST select "Leave Blank" from the drop-down menu for EACH of the remaining "Account" submission boxes (you can leave the "Amount" boxes blank).
  4. For transactions 3, 4, 5, and 8, you are given additional instructions. Read them carefully.
  5. You get 5 tries for each transaction (8 tries for transaction #8).
  6. The entries for each transaction are worth 2 points (4 points for transaction #8).

Transaction 1
On March 1, the three classmates opened a checking account for The Wire at a local bank. They each deposited $22,000 in exchange for shares of stock. A few of their friends also purchased stock totaling $10,000 that was deposited in The Wire account.

(OPTIONS TO ENTER ANSWERS FOR ALL TRANSACTIONS)

Account: Dollar amount:______

Account: Dollar amount:______

Account: Dollar amount:______

Account: Dollar amount:______

Account: Dollar amount:______

Account: Dollar amount:______

Account: Dollar amount:______

Account: Dollar amount:______

Options for account: Cash, accounts receivable, Inventory, Prepaid Rent, Fixtures and Equipment, Accounts Payable, Interest Payable, Wages Payable, Notes Payable, Paid-in Capital, Retained Earnings, Leave in Bank

Transaction 2
The company quickly acquired $42,000 in inventory, 60% of which was paid for in cash. The rest was acquired on open accounts that were payable after 30 days.

Transaction 3
A one-year store rental lease was signed on March 1 for $14,400 for the year, and rent for the first 2 months was paid in advance. [Note: Record the complete entry for the March 1 transaction first and the complete adjusting entry on March 31 second.]

Transaction 4
The owners paid $2,500 for website advertising. They were able to get a good deal because one of the company's owners also owns stock in the website company. The owners also paid $5,500 for some advertising in local newspapers. [Note:Combine both transactions into one entry].

Transaction 5
Sales were $80,000. Cost of merchandise sold was 55% of its sales price. 75% of the sales were on open account. [Note:Record the complete entry for the sales first and the complete entry for the expenses second]

Transaction 6
Wages and salaries in March were $10,000, of which $8,000 was actually paid to employees.

Transaction 7
Miscellaneous expenses were $1,300, all paid for with cash.

Transaction 8
On March 1, fixtures and equipment were purchased for $6,000 with a downpayment of $1,000 and a $5,000 note, payable in one year. Interest of 5.5% per year was due when the note was repaid. The estimated life of the fixtures and equipment is 10 years with no expected salvage value. [Note: Record the complete entry for the March 1 equipment purchase first, the March 31 depreciation adjusting entry second, and the March 31 interest adjusting entry third.  Also, round all answers to the nearest cent.]
Answer #1
1
Account Dollar amount
Cash 76000 =(22000*3)+10000
Paid in capital 76000
2
Account Dollar amount
Cash -25200
Inventory 42000
Accounts Payable 16800
3
Account Dollar amount
Cash -2400
Prepaid rent 2400
Prepaid rent -1200
Retained earnings -1200
4
Account Dollar amount
Cash -8000 =2500+5500
Retained earnings -8000
5
Account Dollar amount
Cash 20000
Accounts Receivable 60000 =80000*75%
Retained earnings 80000
Inventory -44000 =80000*55%
Retained earnings -44000
6
Account Dollar amount
Cash -8000
Wages payable 2000
Retained earnings -10000
7
Account Dollar amount
Cash -1300
Retained earnings -1300
8
Account Dollar amount
Cash -1000
Fixtures and Equipment 6000
Notes Payable 5000
Fixtures and Equipment -50.00 =6000/10/12
Retained earnings -50.00
Interest Payable 22.92 =5000*5.5%/12
Retained earnings -22.92